OpenStudy (anonymous):

Solve the quadratic equation: 2x2 + 11x − 6 = 0

OpenStudy (anonymous):

You can either complete the square or use quadratic formula

OpenStudy (anonymous):

you can even factor if you like

OpenStudy (anonymous):

via \[(x+6)(2x-1)=0\] and go from there

OpenStudy (vishweshshrimali5):

u can use the formula \[x = (-b \pm \sqrt{b^2 - 4ac})/2a\]

OpenStudy (anonymous):

nope not -3,1

OpenStudy (anonymous):

hero come chat with me

OpenStudy (vishweshshrimali5):

now see b = 11 a=2 c=-6 put these values in the formula

OpenStudy (anonymous):

if you factor you get \[x+6=0\] \[x=-6\] or \[2x-1=0\] \[2x=1\] \[x=\frac{1}{2}\]

OpenStudy (vishweshshrimali5):

who is this hero ?

OpenStudy (anonymous):

cherri

OpenStudy (anonymous):

come to twiddla

OpenStudy (vishweshshrimali5):

the values of x = -12 and 1/2

OpenStudy (anonymous):

I'm pretty sure satellite has it

OpenStudy (vishweshshrimali5):

sorry i mean x=-6 or 1/2

OpenStudy (vishweshshrimali5):

yaaa

OpenStudy (vishweshshrimali5):

thats why he is a sensei